OpenStudy (anonymous):

find the derivative of y=(1+cosx)/(1-cosx) Find the derivative of y=4sinx-5cosx+x

OpenStudy (anonymous):

The second function's derivative is pretty straightforward, so I'll leave it to you. For the first, use the quotient rule: \[y'=\frac{(1-\cos x)(-\sin x)-(1+\cos x)(\sin x)}{(1-\cos x)^2}\] Simplify.

OpenStudy (anonymous):

for the first one I simplified and got y'=-2sinx/1-cos^2x I just want to check to see if i did it right

OpenStudy (anonymous):

The numerator is correct, but the denominator is not. \[(1-\cos x)^2\not=1-\cos^2x\]

OpenStudy (anonymous):

thanks I figured it out!

OpenStudy (anonymous):

yw